OpenStudy (anonymous):

[3.07] Choose the best description of how to calculate f[g(−4)] if f(x) = 2x + 1 and g(x) = x2 − 10.

OpenStudy (anonymous):

that was wrong sorry first compute \[g(-4)=(-4)^2-10=16-10=6\] then compute \(f(6)\)

OpenStudy (anonymous):

you got the last step?

OpenStudy (anonymous):

no its difficult can u explain

OpenStudy (anonymous):

first i replaced \(x\) by \(-4\) in the function \(g(x)=x^2-10\)

OpenStudy (anonymous):

just make the replacement, make sure to use parentheses so i got \(g(-4)=(-4)^2-10=(-4)\times (-4)-10=16-10=6\)

OpenStudy (anonymous):

then we we have to replace \(x\) in \(f(x)=2x+1\) by 6 to get \[f(6)=2\times 6-1=12-1=11\] then we are done

OpenStudy (anonymous):

oooh now i get it so much easier when you explain it to me

OpenStudy (anonymous):

hope it helped, i don't think i skipped any steps